# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

### v2

@@ Abstract
Recent advances in large audio language models (LALMs) have greatly enhanced multimodal conversational systems. However, existing benchmarks remain limited—they are mainly English-centric, rely on synthetic speech, and lack comprehensive, discriminative evaluation across multiple dimensions. To address these gaps, we present Voice Chat Bot Bench (VCB Bench)—a high-quality Chinese benchmark built entirely on real human speech. VCB Bench evaluates LALMs from three complementary perspectives: instruction following (including speech-level control beyond text commands), knowledge understanding (general knowledge, reasoning, and daily dialogue), and robustness (stability under perturbations in content, environment, and speaker traits). Experiments on representative LALMs reveal notable performance gaps and highlight future directions for improvement. VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.  [^1] Code and data are available at https://github.com/Tencent/VCB-Bench
@@ 1 Introduction
In recent years, large language models (LLMs) (Vaswani et al., 2017; Anil et al., 2023) have achieved remarkable progress in natural language understanding and generation. Integrating language modeling with modalities such as vision and audio (Radford et al., 2021; Singh et al., 2022) has further given rise to a new paradigm of multimodal learning. Within this trend, large audio language models (LALMs)—which combine speech signal processing with language modeling—have developed rapidly. Emerging systems such as StepAudio2 (Wu et al., 2025) and Qwen3-Omni (Xu et al., 2025b) demonstrate end-to-end (E2E) speech understanding and generation with capabilities in voice question answering, real-time conversation, and audio content analysis. Consequently, voice conversational agents powered by LALMs are drawing increasing academic and industrial attention, offering more natural and human-like interactions than text-only systems.
@@
Despite these advances, moving from basic LALM functionalities to practical voice agents requires reliable and comprehensive evaluation tools. Such benchmarks are essential for diagnosing model weaknesses, guiding optimization, and enabling fair comparisons across systems. While initial efforts (Chen et al., 2024; Yang et al., 2024; Lin et al., 2025) have explored instruction following, audio understanding, reasoning, and dialogue scenarios, current evaluation practices remain limited in three major ways. First, most benchmarks are English-centric, leaving Chinese—the world’s most widely spoken language—largely unexplored. Second, the majority rely on synthetic speech data, which poorly reflects real-world acoustic variability. Third, many are text-derived benchmarks (e.g., AlpacaEval (Li et al., 2023), IFEval (Zhou et al., 2023)), whose formal and lengthy content is unsuitable for evaluating conversationally grounded LALMs that should generate natural, colloquial speech. Addressing these limitations is critical given China’s large user base and the growing demand for practical, high-quality voice agents.
@@
To bridge these gaps, we introduce Voice Chat Bot Bench (VCB Bench)—the first comprehensive evaluation framework for Chinese voice conversation, built entirely from authentic (non-synthetic) speech. VCB Bench evaluates LALMs along three complementary dimensions: (1) Instruction following, extending beyond text-based prompts to incorporate speech-level control tasks such as adjusting volume, speed, and emotion, with bilingual (Chinese-English) support; (2) Knowledge, including multi-disciplinary general knowledge (12 subjects), mathematical and logical reasoning, daily dialogue comprehension, and story continuation for pretraining performance assessment; (3) Robustness, measuring model stability under real-world perturbations across content (mispronunciations, grammatical errors), environment (street, TV noise), and speaker characteristics (age, accents).
@@
Our proposed VCB Bench is built entirely from authentic human recordings rather than synthetic speech. It provides a large-scale, high-fidelity dataset covering diverse conversational scenarios and introduces a multi-dimensional evaluation framework that jointly measures knowledge understanding, instruction following, and robustness through fine-grained, reproducible tasks. Based on this benchmark, we conduct a systematic empirical analysis of state-of-the-art LALMs under unified settings, revealing their strengths and limitations in Chinese voice interaction and offering actionable insights for future model development.
@@ 2 Related Work
Large Audio Language Models. Recent LALMs primarily adopt an E2E audio-language modeling paradigm, integrating speech understanding and generation within a unified framework.
@@
StepAudio models (Huang et al., 2025; Wu et al., 2025) focus on tightly coupling recognition and synthesis. StepAudio integrates a dual-codebook tokenizer and achieves a remarkably low WER, while StepAudio2 advances to a fully E2E design with fixed text-speech token alignment and Chain-of-Thought reasoning, improving fine-grained paralinguistic understanding.
Baichuan-Audio (Li et al., 2025) employs hierarchical RVQ discretization and dual audio heads to balance acoustic and linguistic objectives, enabling real-time bilingual communication. GLM4-Voice (Zeng et al., 2024) introduces a three-module structure (Tokenizer-Backbone-Decoder) supporting emotion and dialect modeling. Kimi-Audio (Ding et al., 2025) fuses continuous acoustic and discrete semantic tokens in a dual-head architecture, achieving low-latency, high-fidelity streaming generation.
These models demonstrate rapid progress in unified audio-language modeling—covering tokenization, multimodal fusion, and real-time dialogue—but systematic benchmarks, especially for Chinese real-speech interaction, remain scarce. Current evaluations are mostly qualitative or based on synthetic data, underscoring the need for comprehensive real-speech benchmarks like VCB Bench.
Audio Benchmarks. Recent efforts have introduced several benchmarks to evaluate LALMs from different perspectives. VoiceBench (Chen et al., 2024) assesses general knowledge, instruction adherence, safety, and robustness, mainly based on existing text datasets such as AlpacaEval and SD-QA (Faisal et al., 2021), with lengthy or highly complex samples removed. OpenAudioBench, released alongside Baichuan-Audio, integrates question-answering datasets including Spoken LLaMA Questions (Nachmani et al., 2023) and Web Questions (Berant et al., 2013), and augments them with a TTS-generated reasoning subset.
AIR Bench (Yang et al., 2024) contains two components—a basic benchmark covering emotion recognition, ASR, and age estimation, and a dialogue benchmark evaluating auditory understanding and internal knowledge. MMAU (Kumar et al., 2025) and MMAR (Ma et al., 2025) focus on deep audio reasoning, requiring multi-step inference grounded in internal audio knowledge. OmniBench (Li et al., 2024) targets omni-modal models handling audio, images, and text, where text queries are paired with multimodal contexts (speech, music, or sound) to test integrated reasoning. Finally, URO Bench (Yan et al., 2025) provides a bilingual (English-Chinese) comprehensive set that evaluates audio understanding, reasoning, and conversational ability—but the speech data are entirely synthetic (TTS-generated).
@@ 3 VCB Bench
As shown in the Figure 1, VCB Bench covers three core dimensions: Instruction Following, Knowledge, and Robustness. Instruction Following includes Text Instruction Follow (TIF) (e.g., continuation, creation), Speech Instruction Follow (SIF) (e.g., emotional, volume control), and Multi-turn Dialog (MTD) tasks. The Knowledge module assesses General Knowledge (GK) across 12 disciplines, Mathematical Logic (ML), Discourse Comprehension (DC), and Story Continuation (SC). Robustness introduces real-world perturbations from speaker variations, environmental noise, and content modifications to evaluate model stability.
@@ 3.1 Dataset Construction
The VCB Bench dataset integrates data from three distinct sources: third-party professional recordings, audio extracted from variety show Q&A segments and an internally curated two-person conversational dialogue dataset. Each source supports different evaluation modules within the benchmark.
Third-Party Recorded Data. This category supports the Instruction Following, ML, and SC tasks under the Knowledge module, as well as the Robustness module. The production pipeline involves the following steps: First, task types and examples are defined through team discussion. Next, commissioned personnel manually compose texts that fulfill the task requirements. These texts then undergo manual inspection to ensure quality. Approved texts are forwarded to a third-party recording team for professional audio production. After recording, the data team performs quality checks on the audio. Subsequently, GPT-4o-Audio is used to evaluate audio quality, while GPT-4o assesses textual quality. Finally, manual screening is conducted to select high-quality samples, determining the final evaluation dataset.
For Robustness data, it’s text materials are derived from Instruction Following module. The original audio from this module serves as the control group. To control for speaker variability, the same speaker re-recorded the text under specified interference conditions (e.g., accent, noisy environment) wherever possible, using the original audio as a baseline. For "content variation" types, the text was first modified (e.g., introducing grammatical or pronunciation errors) before being re-recorded by the same speaker. Additionally, to test performance in extreme scenarios, subsets like Volume, Speed, and Unstable Signal underwent post-processing.
Variety Show Q&A Data. This category supports GK in the Knowledge module. The process includes: crawling about 20 hours of Q&A audio; manual annotation and segmentation for timestamps; ASR transcription; quality scoring (assessing question clarity and answer accuracy) and subject classification, both via GPT-4O; selection of Q&A pairs above a score threshold; and a final manual review for transcription accuracy, answer correctness, and audio clarity.
@@
Internal Two-person Dialogue Dataset. Designed to support the DC module, this category’s data is processed as follows: the original long-form audio undergoes a two-stage segmentation with GPT-4O—first by topic, then refined into semantically coherent segments under one minute. From the transcriptions, GPT-4O generates task-specific QA pairs (e.g., for analysis or induction), followed by a final manual screening to verify question quality and answer accuracy.
@@ 3.2 Dataset Details
Instruction Following. The Instruction Following section comprehensively evaluates LALMs’ ability to understand and execute both text and speech instructions, covering three sub-tasks: TIF, SIF, and MTD. All tasks are open-ended, and both TIF and SIF support Chinese and English to meet cross-lingual evaluation needs.
@@
Model Instruction Following Knowledge Robustness
TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.53 85.05 45.53 62.14 48.64 76.09 80.25 84.41
Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.86 85.66 53.51 79.94 74.76 68.25 75.44 77.71
Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.85 55.43 80.39 73.72 82.31 79.81 80.44
Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.20 44.48 60.33 54.38 65.53 67.62 67.46
Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.89 / / 85.57 35.83 59.28 67.07 83.01 78.11 79.60
StepAudio (Huang et al., 2025) 87.16 66.93 80.25 63.66 / 60.42 77.08 59.52 71.63 74.87 85.81
StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.14 87.72 61.15 81.29 83.08 81.61 77.84 83.04
Mimo-Audio 91.22 91.76 72.89 24.25 / 56.58 84.01 87.92 84.43 86.19 95.10
GPT4o-Audio 91.24 91.66 88.16 86.07 / 61.29 77.68 77.64 82.32 82.49 92.34
Qwen3-Omni (Xu et al., 2025b) 95.59 91.24 84.37 77.50 91.13 74.74 94.27 94.26 93.58 92.68 97.61
@@ 4.1 Configuration
We evaluate the latest and most capable LALMs. The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, GPT4o-Audio and Qwen3-Omni [^2] Qwen3-Omni was included solely in the experimental comparison in Section 4.2..
For the SIF tasks in both Chinese and English (SIF-En), we invoke each model’s “audio2audio” API to generate spoken responses. The adherence to instruction requirements is then automatically scored using GPT-4o-audio. For other tasks except SC, we call the “audio2text” API to obtain textual responses, which are evaluated by GPT-4O. In open-ended question answering, GPT-4O provides a numerical score on a 1-5 scale, while for reference-based QA, it returns a binary "Yes" or "No" judgment.
For the SC task, we assess a subset of pre-trained base models: Baichuan-Audio-Base, Kimi-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base. Following the StoryCloze evaluation protocol, we compute the negative log-likelihood for both the correct and incorrect endings, with model selection determined by comparing these two values. For SIF tasks, the top six performing models undergo further Mean Opinion Score (MOS) evaluation. We sample the first 30 items from each relevant dataset, and eight expert evaluators rate the generated audio samples.
In the MTD evaluation, the model receives input context only in audio. We adopt Bai et al. (2024)’s protocol, requiring the model to answer each dialogue turn using the original ground-truth context, not its prior responses. A key scoring distinction is the heightened focus on the final turn: it carries 50% of the total score per sample, and the first several turns account for the remaining 50%. All Experiments are conducted on H20.
4.2 Main Result
Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the appendix. As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks. GPT-4o-Audio, as a strong non-open-source model, also performs robustly, particularly in SIF.
Figure 4: The investigation of the text-speech alignment capability of LALMs. A2T (Audio-to-Text, generating text directly from audio input), A2A W/ ASR (Audio-to-Audio, transcribing the generated audio via Automatic Speech Recognition to text), and A2A W/O ASR (evaluating the generated audio’s text without ASR). All test sets are from TIF/TIF-En (e.g., “Creat.” for Creation, “Recom.” for Recommendation).
In Instruction Following, Mimo-Audio excels in TIF (91.22) and TIF-En (91.76) with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation. For SIF, GPT-4o-Audio leads in Chinese SIF (88.16) and SIF-En (86.07), while StepAudio series and Kimi-Audio perform robustly in Chinese SIF but their SIF-En scores lag significantly, reflecting challenges in handling English speech’s paralinguistic features. In MTD, StepAudio2Mini (87.72) and Qwen3-Omni (91.13) lead among models, outperforming counterparts like Baichuan-Audio-Chat (73.20), which highlights divergence in long-context dialogue logic control.
In Knowledge, Qwen3-Omni stands out in GK (74.74), ML (94.27), and DC (94.26), surpassing other models and demonstrating strengths in multi-disciplinary knowledge and deep reasoning. Mimo-Audio also performs strongly in ML (84.01) and DC (87.92), suggesting capabilities in deep reasoning and text semantic analysis. However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
Overall, Qwen3-Omni demonstrates comprehensive superiority, while other open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation.

### v3

@@ Abstract
While large audio language models (LALMs) have driven significant progress in multimodal conversational systems, current benchmarks suffer from critical limitations: they are largely English-centric, use synthetic speech, and fail to provide comprehensive, discriminative evaluation across key dimensions. To fill this gap, we present Voice Chat Bot Bench (VCB Bench), a novel, high-quality Chinese benchmark built exclusively on real human speech. VCB Bench assesses LALMs across three complementary axes: instruction following (including speech-level control beyond text commands), knowledge understanding (including general knowledge, reasoning, and daily dialogue), and robustness (evaluating stability under variations in content, environment, and speaker characteristics). Experiments conducted on representative LALMs reveal notable performance disparities and offer tangible insights for future improvements. VCB Bench serves as a reproducible and fine-grained framework, providing standardized evaluation and practical guidance for the development of Chinese voice conversational models.  [^1] Code and data are available at https://github.com/Tencent/VCB-Bench
@@ 1 Introduction
In recent years, large language models (LLMs) (Vaswani et al., 2017; Anil et al., 2023) have achieved remarkable progress in natural language understanding and generation. Integrating language modeling with modalities such as vision and audio (Radford et al., 2021; Singh et al., 2022) has further given rise to a new paradigm of multimodal learning. Within this trend, LALMs—which combine speech signal processing with language modeling—have developed rapidly. Emerging systems such as Step-Audio 2 (Wu et al., 2025) and Qwen3-Omni (Xu et al., 2025b) demonstrate end-to-end (E2E) speech understanding and generation with capabilities in voice question answering, real-time conversation, and audio content analysis. Consequently, voice conversational agents powered by LALMs are drawing increasing academic and industrial attention, offering more natural and human-like interactions than text-only systems.
@@
Despite these advances, the transition from basic LALM functionalities to practical voice agents necessitates reliable and comprehensive evaluation tools. Such benchmarks are essential for diagnosing model weaknesses, guiding optimization efforts, and enabling fair comparisons across different systems. While initial efforts (Chen et al., 2024; Yang et al., 2024; Lin et al., 2025) have explored instruction following, audio understanding, reasoning, and dialogue scenarios, current evaluation practices remain limited in three major aspects: First, existing benchmarks are predominantly English-centric, resulting in insufficient exploration of Chinese, despite it being the world’s most widely spoken language. Second, the majority rely on synthetic speech data, which poorly reflects the acoustic variability of real-world environments. Third, many are derived from text-based benchmarks (e.g., AlpacaEval (Li et al., 2023), IFEval (Zhou et al., 2023)). The formal and lengthy content of these benchmarks is unsuitable for evaluating conversationally grounded LALMs, which are designed to generate natural, colloquial speech. This effort is critical due to China’s vast user base and the rapidly growing demand for practical, high-quality voice agents.
@@
We address these limitations by introducing VCB Bench: the first comprehensive evaluation framework for Chinese voice conversation, built entirely upon authentic (non-synthetic) speech. VCB Bench assesses LALMs across three complementary dimensions: (1) Instruction Following: This extends beyond text-based prompts to include speech-level control tasks, such as adjusting volume, speed, and emotion, with inherent bilingual (Chinese-English) support. (2) Knowledge: This encompasses multi-disciplinary general knowledge (across 12 subjects), mathematical and logical reasoning, daily dialogue comprehension, and story continuation to assess pre-training performance. (3) Robustness: This measures the model’s stability under real-world perturbations, including variations in content (mispronunciations, grammatical errors), environment (street, TV noise), and speaker characteristics (age, accents).
@@
Our proposed VCB Bench is entirely derived from authentic human recordings, moving beyond the limitations of synthetic speech. It provides a large-scale, high-fidelity dataset encompassing diverse conversational scenarios and introduces a multi-dimensional evaluation framework. This framework jointly measures instruction following (including fine-grained control), knowledge understanding, and robustness through a set of reproducible tasks. Based on this benchmark, we conduct a systematic empirical analysis of state-of-the-art LALMs under unified settings, which elucidates their strengths and limitations in Chinese voice interaction and yields actionable insights for future model development.
@@ 2 Related Work
Large Audio Language Models.  LALMs primarily adopt an E2E audio-language modeling paradigm (Tang et al., 2023; Fu et al., 2024; Xie and Wu, 2024), integrating speech understanding and generation within a unified framework.
@@
StepAudio models (Huang et al., 2025; Wu et al., 2025) focus on tightly coupling recognition and synthesis. StepAudio integrates a dual-codebook tokenizer and achieves a remarkably low WER, while Step-Audio 2 mini advances to a fully E2E design with fixed text-speech token alignment and Chain-of-Thought reasoning, improving fine-grained paralinguistic understanding.
Baichuan-Audio (Li et al., 2025) employs hierarchical RVQ discretization and dual audio heads to balance acoustic and linguistic objectives, enabling real-time bilingual communication. GLM-4-Voice (Zeng et al., 2024) introduces a three-module structure (Tokenizer-Backbone-Decoder) supporting emotion and dialect modeling. Kimi-Audio (Ding et al., 2025) fuses continuous acoustic and discrete semantic tokens in a dual-head architecture, achieving low-latency, high-fidelity streaming generation.
In addition, there are many other excellent LALMs, such as MiMo-Audio (Xiaomi, 2025) and Fun-Audio-Chat(Chen et al., 2025). All of these models demonstrate rapid progress in unified audio-language modeling—covering tokenization, multimodal fusion, and real-time dialogue—but systematic benchmarks, especially for Chinese real-speech interaction, remain scarce. Current evaluations are mostly qualitative or based on synthetic data, underscoring the need for comprehensive real-speech benchmarks like VCB Bench.
Audio Benchmarks.  Recent efforts have introduced several benchmarks to evaluate LALMs from different perspectives. VoiceBench (Chen et al., 2024) assesses general knowledge, instruction adherence, safety, and robustness, primarily by adapting content from existing text datasets such as AlpacaEval and SD-QA (Faisal et al., 2021), often after removing lengthy or overly complex samples. OpenAudioBench, released with Baichuan-Audio, integrates question-answering datasets like Spoken LLaMA Questions (Nachmani et al., 2023) and Web Questions (Berant et al., 2013), and further augments them with a TTS-generated reasoning subset.
@@
AIR Bench (Yang et al., 2024) comprises two components: a basic benchmark covering emotion recognition, automatic speech recognition (ASR), and age estimation, and a dialogue benchmark focused on evaluating auditory understanding and internal knowledge. MMAU (Kumar et al., 2025) and MMAR (Ma et al., 2025) concentrate on deep audio reasoning, specifically requiring multi-step inference grounded in the model’s internal audio knowledge. OmniBench (Li et al., 2024) targets omni-modal models that process audio, images, and text. It utilizes text queries paired with multimodal contexts (speech, music, or sound) to test integrated reasoning capabilities. Finally, URO Bench (Yan et al., 2025) provides a comprehensive bilingual (English-Chinese) set for evaluating audio understanding, reasoning, and conversational ability, but its speech data are entirely synthetic (TTS-generated).
@@ 3 VCB Bench
As shown in the Figure 1, VCB Bench covers three core dimensions: Instruction Following, Knowledge, and Robustness. Instruction Following includes Text Instruction Following (TIF) (e.g., continuation, creation), Speech Instruction Following (SIF) (e.g., emotional, volume control), and Multi-turn Dialog (MTD) tasks. The Knowledge module assesses General Knowledge (GK) across 12 disciplines, Mathematical Logic (ML), Discourse Comprehension (DC), and Story Continuation (SC). Robustness introduces real-world perturbations from speaker variations, environmental noise, and content modifications to evaluate model stability. The detailed statistics for each subtask is presented in Appendix 3.
@@ 3.1 Dataset Construction
The VCB Bench dataset integrates data from three distinct sources: third-party professional recordings, audio segments extracted from variety show Q&A, and an internally curated two-person conversational dialogue dataset. Each source is utilized to support different evaluation modules within the benchmark.
Third-Party Recorded Data.  This data category supports the Instruction Following, ML, and SC tasks in the Knowledge module, as well as the Robustness module.The production pipeline has several steps: First, task types and specific examples are defined. Next, professional writers manually craft task-compliant text. This text undergoes rigorous manual quality checks before being sent to a third-party team for professional audio recording. After recording, the data team first checks audio quality to filter out low-quality files. They then conduct detailed manual screening, selecting top-quality samples to form the final evaluation dataset.
The text materials for the Robustness data are derived directly from the Instruction Following module, with the original audio from that module serving as the control group. To control for speaker variability, the same speaker re-recorded the text under specified interference conditions (e.g., accent, noisy environment) wherever possible, using the original audio as a baseline. For content variation types, the text was first modified (e.g., introducing grammatical or pronunciation errors) before being re-recorded by the same speaker. Additionally, to test performance in extreme scenarios, subsets such as Volume, Speed, and Unstable Signal were created using post-processing techniques.
Variety Show Q&A Data.  This category supports GK in the Knowledge module. The process involves: crawling about 20 hours of Q&A audio; manual timestamp annotation and segmentation; ASR transcription; question subject classification; and a final manual review for transcription accuracy, audio clarity, problem statement and answer correctness.
@@
Internal Two-person Dialogue Dataset.  Designed to support the DC module, this category’s data is processed as follows: the original long-form audio undergoes a two-stage segmentation —first by topic, then refined into semantically coherent segments under one minute. From the ASR transcriptions corresponding to the segments, we ask LLM to generate task-specific QA pairs (e.g., for analysis or induction), followed by a final manual screening to verify question quality and answer accuracy.
@@ 3.2 Dataset Details
Instruction Following.  The Instruction Following section provides a comprehensive evaluation of LALMs’ capacity to understand and execute both text and speech commands, encompassing three sub-tasks: TIF, SIF, and MTD. All tasks are open-ended, and both TIF and SIF support bilingual evaluation (Chinese and English) to address cross-lingual assessment requirements.
@@
Model Instruction Following Knowledge Robustness
TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
GLM-4-Voice (Zeng et al., 2024) 82.15 75.52 73.18 66.94 82.56 41.79 60.18 / 73.64 77.51 78.60
Kimi-Audio (Ding et al., 2025) 77.33 60.37 71.04 54.13 82.27 54.47 75.42 51.96 64.35 65.14 71.51
Qwen2.5-Omni (Xu et al., 2025a) 76.75 60.83 54.04 42.18 85.23 48.13 73.00 72.81 77.82 74.41 78.27
Baichuan-Audio (Li et al., 2025) 82.24 78.60 53.67 48.27 80.96 40.73 74.05 51.66 78.05 77.62 78.71
Step-Audio 2 mini (Wu et al., 2025) 76.10 70.47 60.80 49.85 82.31 49.86 72.85 79.46 75.19 70.58 74.78
MiMo-Audio (Xiaomi, 2025) 90.08 81.89 56.26 42.74 86.30 48.70 81.75 / 83.72 85.36 89.38
GPT-4o-Audio (OpenAI, 2024) 86.94 88.80 77.98 82.90 33.59 55.81 73.45 76.74 80.34 79.92 86.51
Qwen3-Omni (Xu et al., 2025b) 90.45 85.17 70.73 65.57 87.17 66.86 81.90 82.78 87.91 85.63 86.03
Fun-Audio-Chat (Chen et al., 2025) 89.30 89.39 78.82 75.70 / 46.30 73.00 79.15 88.60 83.83 85.15
@@ 4.1 Configuration
We evaluate the latest and most capable LALMs. The selected models comprise GLM-4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio, Step-Audio 2 mini, MiMo-Audio, GPT-4o-Audio, Qwen3-Omni and Fun-Audio-Chat.
For all tasks except SC, we generate spoken responses by calling each model’s audio-to-audio API. For SIF tasks in both Chinese and English, the audio responses are evaluated directly using Gemini-2.5-Pro. For the remaining tasks, following prior work (Yan et al., 2025), we first transcribe the audio responses using an ASR model (Whisper (Radford et al., 2023) for English tasks and Paraformer (Gao et al., 2022) for Chinese tasks) and then assess the transcriptions with GPT-4o. In open-ended QA, both Gemini and GPT assign a numerical score on a 1–5 scale, whereas for reference-based QA, they output a binary judgment of "Yes" or "No."
For the SC task, we assess a subset of pre-trained base models: Baichuan-Audio-Base, Kimi-Audio-Base, Qwen2-Audio-Base, and Step-Audio 2 mini-Base. Following the StoryCloze evaluation protocol, we compute the negative log-likelihood for both the correct and incorrect endings, with model selection determined by comparing these two values. For SIF tasks, six performing models undergo further Mean Opinion Score (MOS) evaluation. We sample the first 30 items from each relevant dataset, and eight expert evaluators rate the generated audio samples.
In the MTD evaluation, we adopt Bai et al. (2024)’s protocol, requiring the model to answer each dialogue turn using the original ground-truth context, not its prior responses. A key scoring distinction is the heightened focus on the final turn: it carries 50% of the total score per sample, and the first several turns account for the remaining 50%. All Experiments are conducted on H20.
@@
4.2 Main Results
Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the Appendix. As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks.
In Instruction Following, Qwen3-Omni (90.45) and MiMo-Audio (90.08) excel in TIF, while Fun-Audio-Chat leads in TIF-En (89.39), closely trailing GPT-4o-Audio (88.80). For SIF, Fun-Audio-Chat (78.82) and GPT-4o-Audio (77.98) lead; GPT-4o-Audio dominates SIF-En (82.90), with most open-source models like Step-Audio 2 mini (49.85) lagging, reflecting challenges in English speech paralinguistic features. In MTD, Qwen3-Omni (87.17) and MiMo-Audio (86.30) lead, while GPT-4o-Audio performs poorly (33.59), highlighting divergence in long-context dialogue control.
In Knowledge, Qwen3-Omni stands out with top scores in GK (66.86), ML (81.90) and DC (82.78), showing superior multi-disciplinary knowledge and reasoning. MiMo-Audio performs strongly in ML (81.75), comparable to Qwen3-Omni. However, Baichuan-Audio (GK:40.73, ML:74.05), GLM-4-Voice (GK:41.79, ML:60.18) and other models lag, revealing gaps in knowledge coverage and step-by-step reasoning.
Overall, Qwen3-Omni demonstrates comprehensive superiority. Other models have task-specific strengths: Fun-Audio-Chat in cross-lingual TIF and SIF, MiMo-Audio in TIF and ML. Most open-source E2E LALMs face English adaptation challenges, while GPT-4o-Audio lacks long-context dialogue capability.